# DART[Uncaptioned image]: An AIGT Detector using AMR of Rephrased Text

### v1

@@ Abstract
As large language models (LLMs) generate more human-like texts, concerns about the side effects of AI-generated texts (AIGT) have grown. So, researchers have developed methods for detecting AIGT. However, two challenges remain. First, the performance on detecting black-box LLMs is low, because existing models have focused on probabilistic features. Second, most AIGT detectors have been tested on a single-candidate setting, which assumes that we know the origin of an AIGT and which may deviate from the real-world scenario. To resolve these challenges, we propose DART[Uncaptioned image], which consists of four steps: rephrasing, semantic parsing, scoring, and multiclass classification. We conducted several experiments to test the performance of DART[Uncaptioned image] by following previous work. The experimental result shows that DART[Uncaptioned image] can discriminate multiple black-box LLMs without probabilistic features and the origin of AIGT.
@@
Hyeonchu Park*, Byungjun Kimthanks: Equal Contribution, and Bugeun Kim Department of Artificial Intelligence, Chung-Ang University, Republic of Korea {phchu0429, k36769, bgnkim}@cau.ac.kr
@@ 1 Introduction
To address these challenges, we propose a Detector using AMR of Rephrased Text (DART[Uncaptioned image]). DART utilizes the semantic gap between given input and rephrased text, using Abstract Meaning Representation (AMR). We assess DART on a single-candidate and a multi-candidate setting to verify the real-world detection performance. As a result, DART can successfully discriminate humans from four cutting-edge LLMs, including GPT-3.5-Turbo, GPT-4o, Llama 3 Dubey et al. (2024), and Gemini-1.5-Flash.
@@
  • We present a semantics-based detection framework for AIGT, leveraging semantic gaps between given text and rephrased texts.
  • DART can discriminate different LLMs and outperform other detectors; e.g., DART achieved a 96.5% F1 score on average.
@@ 2 Background
As a single-candidate performance is far from real-world scenarios, recent AIGT detectors were designed to detect multiple candidates (Li et al., 2023; Abburi et al., 2023; Wang et al., 2023; Shi et al., 2024; Antoun et al., 2024). For example, POGER (Shi et al., 2024) extends resampling methods to estimate probability using about 100 paraphrases. Because of such an excessive regeneration, POGER incurs high computational costs. Besides, SeqXGPT (Wang et al., 2023) used a Transformer-based detector with a proxy model estimating probabilistic features. However, these studies mainly focused on (1) surface-level features and (2) the limited range of LLMs (e.g., GPT family), raising questions about detecting other cutting-edge LLMs.
@@ 3 The DART Framework
DART uses a rephraser module to generate semantically similar text with an LLM. We assume rephrasing can hint at different origins because LLMs and humans may have different writing styles. Mathematically, let us denote the given text as $T_{0}$. Then, we can obtain $T_{1}$ by applying the rephraser on $T_{0}$. We want to utilize a semantic gap between $T_{0}$ and $T_{1}$ to distinguish the origin of AIGT. We additionally used a larger gap between $T_{0}$ and $T_{2}$, a rephrase of $T_{1}$. For the rephraser, we adopted GPT-4o-20240513, as the model showed the highest performance related to semantics-related tasks (OpenAI, 2024a). For the prompts used in the rephraser, see Appendix A.1.
@@
DART uses metrics for semantic parsers to measure semantic gaps between texts. As we adopted AMR as a semantic representation in the previous step, we utilize a fast and efficient algorithm for scoring AMR similarity called SEMA (Anchiêta et al., 2019; Ki et al., 2024). To obtain semantic gaps between $A_{0}$ and $A_{i}$ ($i>0$), DART computes precision $p_{i}$ and recall $r_{i}$ scores generated by SEMA, resulting a feature vector $v=[p_{0},p_{1},r_{0},r_{1}]^{\top}$ for the next step.
@@
DART has a classifier that predicts one possible origin of $T_{0}$. DART uses interpretable classifiers, including support vector machine (SVM) or the decision tree (DT), though one can use any classifier that maps $v$ to one of the candidate origins.
@@ 4 Experiments
To evaluate the performance of DART, we conducted two experiments: (1) single-candidate and (2) multi-candidate. First, in the single-candidate experiment, we adopted a binary classifier as DART’s classifier, assuming that AIGTs are exclusively produced by a specific LLM. Second, in the multi-candidate experiment, we utilized a multiclass classifier, assuming that we were unaware of the precise origin of the AIGT. We ran each experiment 10 times to ensure their reproducibility. Additionally, we analyzed DART’s training efficiency when decreasing the size of training set.
@@ 4.2 Baselines
As baselines, we used five open-source state-of-the-art detectors: DetectGPT Mitchell et al. (2023), Fast-DetectGPT Bao et al. (2024), DNA-GPT Yang et al. (2023), Roberta-base Solaiman et al. (2019), and SeqXGPT Wang et al. (2023). Among these models, DetectGPT, Fast-DetectGPT, and SeqXGPT used probabilistic features generated by third-party LLMs in order to detect cutting-edge LLMs. Meanwhile, DNA-GPT and Roberta-base used shallow semantic features, such as N-grams or contextual embeddings. Note that DART stands out from these models because it uses AMR-based semantic features rather than probabilistic features.
We used a different set of models for the two experiments, considering experiments reported with five models. For the single-candidate experiment, we compared DART with all five models. For the multi-candidate experiment, we compared DART only with SeqXGPT, as it is the only model tested on the multi-candidate setting. To ensure a fair comparison, all models used in the experiment are trained on our dataset from scratch [^1] Note that we used GPT-2 as a proxy model for the GPT series and Gemini for the models using probabilistic features because they do not provide log probability via API, following Bao et al. (2024).. To measure the performance, we used the F1 score.
GPT series Llama Gemini
Avg 3.5T 4o 3 1.5F
DetectGPT* 65.8 65.8 65.6 65.8 65.7
f-DetectGPT* 60.1 58.0 66.2 62.4 53.8
DNA-GPT 54.1 56.6 57.4 54.8 47.7
Roberta-base 77.2 76.8 80.0 74.7 77.1
SeqXGPT* 54.1 86.5 45.9 41.6 42.3
DARTSVM 82.8 87.1 86.1 84.8 73.3
   DARTDT 96.5 100.0 88.1 100.0 97.9
@@
Table 1: Results of the single-candidate experiment
@@ 5 Result and Discussion
Mac. GPT series Llama Gemini Human
F1 3.5T 4o 3 1.5F
SeqXGPT* 59.2 54.3 66.4 44.8 61.4 69.3
DARTSVM 65.0 67.4 71.0 54.0 67.0 65.4
  DARTDT 81.2 80.6 76.6 85.8 85.5 77.3
@@
Table 2: Results of the multi-candidate experiment
Single-candidate experiment
DART outperformed existing models. As shown in Table 1, DARTDT and DARTSVM achieved 96.5% and 82.8% F1 scores on average, which are 19.3%p and 5.6%p higher than the Roberta-base model (77.2%). Also, DARTDT can detect all four cutting-edge models with over 85% of F1 score. Meanwhile, other existing models showed F1 scores lower than 70%, on average. Their F1 scores are sometimes lower than the random binary baseline (50%): DNA-GPT on detecting Gemini and SeqXGPT on detecting LLaMA and Gemini.
@@
We suspect that DARTDT can achieve such outstanding performance because DART’s semantic scoring step can successfully form several clusters according to their origins. To support this argument, we further analyzed the feature vectors of DART using principal component analysis. We found that several clusters are formed for each origin, as described in Appendix C.1.
Multi-candidate experiment
@@
DART also showed much higher performance than SeqXGPT. As shown in Table 2, DARTDT and DARTSVM achieved 81.2% and 65.0% macro F1 scores, which are 22.0%p and 5.8%p higher than SeqXGPT (59.2%). Interestingly, SeqXGPT achieved the lowest F1 score on Llama 3 (44.8%), but DARTDT achieved the lowest score on GPT-4o (76.6%).
Figure 2: Contingency matrix from multi-candidate experiment. Left (a) and Right (b) correspond to SeqXGPT and DARTDT. Actual and predicted classes are depicted as horizontal and vertical axes.
@@
We suspect how the detectors extract features using an LLM affects the performance. To support this claim, we present a contingency table of SeqXGPT and DARTDT, as shown in Figure 2. The figure shows that (i) SeqXGPT struggled in distinguishing models other than Llama 3, and (ii) DARTDT struggled in distinguishing the GPT family and humans. Since SeqXGPT in our experiment used GPT-2 as a proxy model DARTDT used GPT-4o as a rephraser module, the characteristics of such LLMs affected the detection performances. For example, as DARTDT utilizes semantic gaps between the original text and rephrased texts, origins should reveal distinguishable gaps to successfully identify them. That is, when the gaps are too similar between origins to discriminate them, DARTDT faces difficulty in the classification step. Since GPT-4o has a similar language understanding ability to humans (OpenAI, 2024a), GPT-4o and humans may be less distinguishable through gaps. Similarly, as GPT-3.5-turbo may share some core knowledge with GPT-4o, GPT-4o can be confused with GPT-3.5T in DARTDT.
@@
Figure 3: Training efficiency on the multi-candidate experiment
Training efficiency of DART
@@
Figure 3 shows the result of the experiment on training efficiency. The result shows that DARTDT is robust even though we use a small amount of training data. Specifically, DARTDT maintained the F1 score until we used 1/32 of the training set (about 500 examples). Meanwhile, the performance of SeqXGPT and DARTSVM monotonically decreases as we reduce the size of the training set.
@@ 6 Conclusion
We introduced an AIGT framework, DART[Uncaptioned image]  to tackle challenges in applying AIGT detectors to real-world scenarios. To address the challenges of black-box models, DART employed rephraser and semantic gap scoring module. And, to evaluate whether DART can address vagueness of origin, we assessed DART in two experimental settings: single and multi-candidate settings. As a result, DART achieved outstanding performance compared to existing models, demonstrating successful capture of differences across origins with semantic gaps.
@@
Despite the outstanding performance of DART, this paper has three limitations. First, we tested DART only with a single rephraser LLM, GPT-4o. Though GPT-4o provided enough semantic information to distinguish AIGTs successfully, it is yet questionable whether DART can be used with other rephraser LLMs, such as Llama 3, Gemini Pro, or others. Since different language models may provide different rephrased texts, we need further study to figure out to what extent rephraser LLM affects the performance.
@@
Third, DART tested on a narrow range of black-box models. While narrow LLMs have become publicly available through paid APIs or pretrained parameters. We tried our best to include recent LLMs, such as Gemini Pro or Claude 3. However, we finally excluded those models because their safeguards prevented to generate AIGTs based on a given human-written text when preparing the AIGT dataset. To generalize our findings to other origins, we need a further study on a wider range of models by designing a new way of generating AIGTs.

### v2

@@ Abstract
As large language models (LLMs) generate more human-like texts, concerns about the side effects of AI-generated texts (AIGT) have grown. So, researchers have developed methods for detecting AIGT. However, two challenges remain. First, the performance of detecting black-box LLMs is low because existing models focus on probabilistic features. Second, most AIGT detectors have been tested on a single-candidate setting, which assumes that we know the origin of an AIGT and which may deviate from the real-world scenario. To resolve these challenges, we propose DART[Uncaptioned image], which consists of four steps: rephrasing, semantic parsing, scoring, and multiclass classification. We conducted three experiments to test the performance of DART. The experimental result shows that DART can discriminate multiple black-box LLMs without probabilistic features and the origin of AIGT.
@@
Hyeonchu Park, Byungjun Kim, and Bugeun Kim Department of Artificial Intelligence, Chung-Ang University, Republic of Korea {phchu0429, k36769, bgnkim}@cau.ac.kr
footnotetext: Equal contribution.
@@ 1 Introduction
To address these challenges, we propose a Detector using AMR of Rephrased Text (DART[Uncaptioned image]). DART utilizes the semantic gap between given input and rephrased text, using Abstract Meaning Representation (AMR). This rephrasing idea was first introduced by RAIDAR (Mao et al., 2024); we adopted a similar idea to reveal such a semantic gap. To examine the real-world detection performance, we assess DART in three settings: single-candidate, multi-candidate, and leave-one-out. Experimental results show that DART can successfully discriminate humans from four cutting-edge LLMs, including GPT-3.5-turbo, GPT-4o, Llama 3-70b Dubey et al. (2024), and Gemini-1.5-Flash.
@@
  • We present a semantics-based detection framework for AIGT, leveraging semantic gaps between given input text and rephrased texts.
  • DART can discriminate different LLMs and outperform other models. On average, DART beat others by more than 19% in F1 score.
  • Also, DART can generalize its knowledge on detecting unseen source models. Specifically, DART achieved a 85.6% F1 score on leave-one-out experiment.
@@ 2 Background
As a single-candidate performance is far from real-world scenarios, recent AIGT detectors were designed to detect multiple candidates (Li et al., 2023; Abburi et al., 2023; Wang et al., 2023; Shi et al., 2024; Antoun et al., 2024). For example, POGER (Shi et al., 2024) extends resampling methods to estimate probability using about 100 paraphrases. Because of such an excessive regeneration, POGER incurs high computational costs. Besides, SeqXGPT (Wang et al., 2023) used a Transformer-based detector with a proxy model estimating probabilistic features. However, these studies mainly focused on surface-level features and the limited range of LLMs (e.g., GPT family), raising questions about detecting other cutting-edge LLMs.
@@ 3 The DART Framework
We hypothesized that rephrasing texts could reveal the difference between humans and AI in the way they express semantics. To obtain the rephrased texts, DART uses a rephraser module that generates semantically closer text $T_{1}$ from a given text $T_{0}$. Further, we let the rephraser generate another rephrased text $T_{2}$ by giving $T_{1}$ to attain additional features. To avoid generating rephrased texts irrelevant to the given input, we need a reliable rephraser that can preserve semantics. So, we adopted GPT-4o-20240513 as our rephraser because the model showed the highest performance in semantics-related tasks OpenAI (2024a). Appendix A.2 details the prompts used in the rephrasing step.
@@
DART uses metrics for semantic parsers to measure semantic gaps between texts. As we adopted AMR as a semantic representation in the previous step, we utilize a fast and efficient algorithm for scoring AMR similarity called SEMA (Anchiêta et al., 2019; Ki et al., 2024). To obtain semantic gaps between $A_{0}$ and $A_{i}$ ($i>0$), DART computes precision $p_{i}$ and recall $r_{i}$ scores generated by SEMA, resulting a feature vector $v=[p_{1},p_{2},r_{1},r_{2}]^{\top}$ for the next step.
@@
DART has a classifier that predicts one possible origin of $T_{0}$. DART uses interpretable classifiers, including support vector machine (SVM) or decision tree (DT), though any classifier that maps $v$ to origins can be used.
@@ 4 Experiments
To evaluate the performance of DART, we conducted three experiments: (1) single-candidate, (2) multi-candidate, and (3) leave-one-out settings. First, in the single-candidate setting, we formulate AIGT detection as a binary classification task. Assuming that AIGTs are exclusively produced by a specific LLM, a detector should predict whether the given text is produced by the LLM. Second, in the multi-candidate setting, we formulate the task as a multi-label classification. After training on AIGTs from multiple candidate sources, a detector should decide the source of the given input text among the candidates. Third, in the leave-one-out setting, we test the generalizability of detectors. We examined whether a detector can successfully classify AIGTs from models that were unseen during the training.
We ran each experiment 10 times for each experiment to achieve reproducibility. Further, we analyzed DART’s training efficiency by examining the decreasing rate of detecting performance as the size of the training dataset.
@@ 4.2 Baselines
As baselines, we used five open-source state-of-the-art detectors: DetectGPT Mitchell et al. (2023), Fast-DetectGPT Bao et al. (2024), DNA-GPT Yang et al. (2023), Roberta-base Solaiman et al. (2019), and SeqXGPT Wang et al. (2023). Among these models, DetectGPT, Fast-DetectGPT, and SeqXGPT used probabilistic features generated by third-party LLMs in order to detect cutting-edge LLMs. Meanwhile, DNA-GPT and Roberta-base used shallow semantic features, such as $n$-grams or contextual embeddings. DART stands out from these models because it uses AMR-based semantics rather than probabilistic features.
We used a different set of detectors for the three experiments, considering experiments reported with five baselines. For the single-candidate experiment, we compared DART with all five detectors. For the multi-candidate and the leave-one-out experiments, we compared DART only with SeqXGPT, as it is the only existing detector that can trained on multiple candidates simultaneously. To ensure a fair comparison, all detectors used in the experiment are trained on our dataset from scratch [^1] Note that we used GPT-2 as a proxy model for the GPT series and Gemini-1.5 when the detectors require probabilistic features because GPT and Gemini do not provide logits, following Bao et al. (2024).. To measure the performance, we used the F1 score.
Average GPT-3.5-turbo GPT-4o Llama3-70B Gemini-1.5
DetectGPT* 65.8 65.8$\pm$ 0.20 65.6$\pm$ 0.16 65.8$\pm$ 0.17 65.7$\pm$ 1.12
fast-DetectGPT* 60.1 58.0$\pm$ 1.94 66.2$\pm$ 0.25 62.4$\pm$ 0.48 53.8$\pm$ 0.58
DNA-GPT 54.1 56.6$\pm$ 1.49 57.4$\pm$ 0.50 54.8$\pm$ 2.60 47.7$\pm$ 2.36
Roberta-base 77.2 76.8$\pm$ 3.24 80.0$\pm$ 2.81 74.7$\pm$ 1.77 77.1$\pm$ 2.13
SeqXGPT* 54.1 86.5$\pm$ 0.48 45.9$\pm$ 0.23 41.6$\pm$ 0.31 42.3$\pm$ 0.52
DARTSVM 82.8 87.1$\pm$ 0.65 86.1$\pm$ 0.70 84.8$\pm$ 2.20 73.3$\pm$ 0.76
  DARTDT 96.5 100.0$\pm$ 0.03 88.1$\pm$ 0.98 100.0$\pm$ 0.03 97.9$\pm$ 1.65
@@
Table 1: F1 scores of detectors in the single-candidate experiment, with standard deviations reported.
@@ 5 Result and Discussion
Single-candidate experiment:
DART outperformed existing models. As shown in Table 1, our DARTDT and DARTSVM achieved 96.5% and 82.8% F1 scores on average, which are 19.3%p and 5.6%p higher than the Roberta-base model (77.2%). Also, DARTDT can detect all four cutting-edge models with over 85% of F1 score. Meanwhile, other existing models showed F1 scores lower than 70%, on average. Moreover, DNA-GPT and SeqXGPT sometimes showed F1 scores lower than the random binary baseline (50%).
Macro F1 GPT-3.5-turbo GPT-4o Llama3-70B Gemini-1.5 Human
SeqXGPT* 59.2$\pm$0.66 54.3$\pm$1.44 66.4$\pm$1.08 44.8$\pm$0.95 61.4$\pm$1.68 69.3$\pm$0.93
DARTSVM 65.0$\pm$0.77 67.4$\pm$0.81 71.0$\pm$1.20 54.0$\pm$1.26 67.0$\pm$0.94 65.4$\pm$1.16
  DARTDT 81.2$\pm$1.71 80.6$\pm$4.61 76.6$\pm$1.16 85.8$\pm$5.42 85.5$\pm$2.36 77.3$\pm$0.88
@@
Table 2: F1 scores of detectors in the multi-candidate experiment, with standard deviations reported.
Macro F1 GPT-3.5-turbo GPT-4o Llama3-70B Gemini-1.5
SeqXGPT* 78.5$\pm$1.04 79.9$\pm$1.39 80.2$\pm$0.52 78.8$\pm$0.92 75.1$\pm$1.31
DARTSVM 56.3$\pm$0.96 56.0$\pm$1.31 59.2$\pm$1.01 56.4$\pm$0.82 53.6$\pm$0.70
  DARTDT 84.2$\pm$1.39 99.3$\pm$0.16 75.8$\pm$3.82 99.1$\pm$0.55 62.5$\pm$1.03
* Models used GPT-2 as a proxy model for black-box models, except Llama3
Table 3: F1 scores of detectors in the leave-one-out experiment, with standard deviations reported.
@@
We suspect that DARTDT can achieve such outstanding performance because our semantic scoring step can successfully form several clusters according to their origins. To support this argument, we further analyzed the feature vectors of DART using principal component analysis. We found that texts sharing the same source usually form several independent clusters rather than spread over the space. Detailed results are presented in Appendix C.3.
Multi-candidate experiment:
@@
DART also outperformed SeqXGPT. As shown in Table 2, our DARTDT and DARTSVM achieved 81.2% and 65.0% macro F1 scores, which are 22.0%p and 5.8%p higher than SeqXGPT (59.2%). Interestingly, SeqXGPT achieved the lowest F1 score on detecting Llama 3 (44.8%), but DARTDT achieved the lowest score on detecting GPT-4o (76.6%).
@@
We suspect how the detectors extract features using an LLM affects the performance. We present a contingency table of SeqXGPT and DARTDT to support this claim, as shown in Figure 2. The figure shows that (i) SeqXGPT struggled in distinguishing models other than Llama 3, and (ii) DARTDT struggled in distinguishing the GPT family and humans. Since SeqXGPT in our experiment used GPT-2 as a proxy model, and DARTDT used GPT-4o as a rephraser module, the characteristics of the used LLMs affected the detection performances. For example, as DARTDT utilizes semantic gaps between the original and rephrased texts, origins should reveal distinguishable gaps to identify them successfully. So, when the gaps are too similar between origins to discriminate them, DARTDT faces difficulty in the classification step.
Figure 2: Contingency matrix from multi-candidate experiment. Top (a) and Bottom (b) correspond to SeqXGPT and DARTDT. Actual and predicted classes are depicted as horizontal and vertical axes.
Since GPT-4o has a similar language understanding ability to humans (OpenAI, 2024a), GPT-4o and humans may be less distinguishable through gaps. Similarly, as GPT-3.5-turbo may share some core knowledge with GPT-4o, GPT-4o can be confused with GPT-3.5-turbo in DARTDT.
Leave-One-Out experiment:
DARTDT showed the best performance. As shown in Table 3, DARTDT achieved 85.6% average F1 score, followed by SeqXGPT (77.9%) and DARTSVM (56.5%). Besides, DARTDT scored 62.5% F1 on detecting the unseen Gemini-1.5, though DARTDT recorded more than 75% on detecting others.
This result indicates that DARTDT can generalize trained knowledge to detect unseen source models. That is, DARTDT can discriminate new candidate models from humans. Specifically, compared to the single-candidate result (Table 1), our model showed almost similar performance on detecting GPT-3.5-turbo and Llama 3 without training on those models. As in the single-candidate experiment, we believe that our semantic scoring step helped to detect unseen models because they form clusters independent from humans. Also, when the cluster becomes indiscernible with humans, DARTDT struggles to detect new models. For example, DARTDT showed a big performance drop when excluding Gemini-1.5 from the training set because DARTDT often confused Gemini-1.5 with humans (top-right corner on Figure 2b).
@@
Figure 3: F1 score of detectors when we decrease the amount of training data in multi-candidate experiment.
Training efficiency of DART:
@@
Here, we discuss the general tendency of the result. Figure 3 shows the performance changes when we decrease the size of the training set. Detailed result of training efficiency is presented in Appendix C.4. The result shows that DARTDT is robust even though we use a small amount of training data. Specifically, DARTDT maintained a similar F1 score until we used 1/32 of the training set (about 500 examples). Meanwhile, the performance of SeqXGPT and DARTSVM monotonically decreases as we reduce the size of the training set.
@@ 6 Conclusion
We introduced an AIGT framework, DART[Uncaptioned image]  to tackle challenges in applying AIGT detectors to real-world scenarios. DART employed rephraser and semantic gap scoring module to address the challenges of black-box models. To evaluate whether DART can address vagueness of origin, we assessed DART in three experimental settings: single-candidate, multi-candidate, and leave-one-out settings. As a result, DART achieved outstanding performance compared to existing AIGT detectors, demonstrating successful capture of differences across origins with semantic gaps.
@@
Despite the outstanding performance of DART, this paper has three limitations. First, we tested DART only with a single rephraser LLM, GPT-4o. Though GPT-4o provided enough semantic information to distinguish AIGTs successfully, it is questionable whether DART can be used with other rephraser LLMs, such as Llama 3, Gemini Pro, or others. Also, we recognize the cost implications of utilizing GPT-4o as a rephraser, which could restrict its applicability in resource-limited environments. Since different language models may provide different rephrased texts with lower costs, we need further study to determine how much rephraser LLM affects the performance.
@@
Third, DART tested on a narrow range of black-box models. While narrow LLMs have become publicly available through paid APIs or pretrained parameters, we tried our best to include recent LLMs, such as Gemini Pro or Claude 3. However, we finally excluded those models because their safeguards prevented from generating AIGTs based on a given human-written text when preparing the AIGT dataset. To generalize our findings to other origins, we need to conduct further studies in a broader range of models and design a new method of generating AIGTs.
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-01341, Artificial Intelligence Graduate School Program, Chung-Ang University)